OpenStudy (anonymous):

Elena bicycles 4 km/h faster than Robert. In the same time it takes Robert to bicycle 48 km, Elena can bicycle 60 km. How fast does each bicyclist travel? Robert travels ? km/h. Elena travels ? km/h.

OpenStudy (anonymous):

let x be elena's speed and y robert's x=y+4 48/y=60/x==>y=(4/5)*x ==>x=(4/5)*x+4===>x/5=4==>x=20 y=16

OpenStudy (anonymous):

THX